OpenStudy (anonymous):

write 35/6 in the improper fraction as a mixed number in simplest form

OpenStudy (anonymous):

\[\frac{ 35 }{ 6 }= \frac{ (6\times5)+5 }{ 6 } = 5 + \frac{ 5 }{ 6 }\] Here you are.

OpenStudy (pixiedust1):

How many times can 6 go into 35?

OpenStudy (anonymous):

5 times. A remainder of 5 is still left.

OpenStudy (pixiedust1):

Correct! So, that means its 5 and then we have a remainder of 5 \[5\frac{ 5 }{ 6}\]

OpenStudy (anonymous):

Yes. But \[5\frac{ 5 }{6}\] is childish. It's actually \[5 + \frac{ 5 }{ 6 }\] because the former can be confused with \[5 \times \frac{ 5 }{ }\]

OpenStudy (anonymous):

I meant confused with \[5 \times \frac{ 5 }{ 6 }\]